# Critical properties of metallic and deconfined quantum critical points in Dirac systems

### v1

@@ I Introduction
In this work, we study a lattice model that features two metallic quantum phase transitions as function of coupling strength [32]. The first one is continuous and can be understood within the $(2+1)$-dimensional Gross-Neveu-SO(3) [GN-SO(3)] field theory [33]. In the SO(3)-broken phase , two out of three Dirac cones aquire a mass, and the wave function of the ungapped Dirac electron couples to the SO(3) order parameter. Similar mass terms have been put forward in spin-orbit coupled fermions on a honeycomb lattice [34]. The second one is a transition between two long-range-ordered phases that break different symmetries: the aforementioned SO(3)-spin-symmetry-broken semimetal and a U(1)-charge-symmetry-broken fully gapped state. The latter can be understood as an insulator characterized by U(1) interlayer coherence [32]. We show, nevertheless, that the interlayer-coherent insulator is degenerate with an $s$-wave superconducting state as a consequence of a partial particle-hole (PPH) symmetry. At the level of mean-field theory, the transition between the SO(3)-spin- and U(1)-charge-symmetry-broken phases is strongly first order [32]. Quantum fluctuations beyond mean-field theory significantly weaken this order-to-order transition, rendering it a candidate for a deconfined quantum critical point in the presence of gapless fermionic excitations—a potential microscopic realization of metallic deconfined quantum criticality [35]. We use large-scale sign-problem free quantum Monte Carlo (QMC) simulations with a Hermitian Trotter decomposition, which significantly improves convergence properties with respect to the limit of small Trotter time steps, to compute the quantum critical and thermodynamic properties of both transitions. Furthermore, in contrast to our initial exploratory study of this model [32], we employ a microscopic implementation that preserves the model’s PPH symmetry explicitly, maintaining the degeneracy of the interlayer-coherent insulating and $s$-wave superconducting ground states in the fully gapped phase already on finite lattices. The interlayer-coherent insulating and $s$-wave superconducting states allow finite-temperature phases above the zero-temperature order. The latter breaks both U(1)-charge and PPH symmetries. Upon increasing temperature, it is therefore possible that the U(1) and PPH orders melt at different temperatures, with an intermediate vestigial phase at intermediate temperatures [36, 37]. We compute the finite-temperature phase diagram of the model and show that the boundaries of the low-temperature orders vanish upon approaching the putative deconfined quantum critical point, in agreement with the expectation for a continuous or weak-first-order transition.
@@ II Model
Figure 1: (a) Sketch of bilayer honeycomb model defined by the Hamiltonian given in Eqs. (1)-(3). Each lattice site (purple dots) contain three SO(3) spin flavor degrees of freedom. The two layers interact via an SO(3)-symmetric spin-density interaction parametrized by the coupling $J$ (orange link). (c) Bottom: Schematic quantum phase diagram as function of $J$, showing the disordered Dirac semimetal (DSM) at small $J$, the SO(3)-spin-symmetry-broken Dirac semimetal (SO3 DSM), in which $2/3$ of the Dirac cones acquire a mass gap, while $1/3$ remains gapless, at intermediate $J$, as well as the fully gapped degenerate interlayer-coherent insulator (ILC) and $s$-wave superconductor (SSC) at strong $J$. The insets illustrate the fermion band structures in the three different phases. Top: Fermion single-particle spectral function from lattice mean-field theory along a high-symmetry path in the first Brillouin zone (as shown in the (b)), illustrating the partial and full gap openings in the SO3 DSM and ILC/SSC phases, respectively.
@@ III.2 SO(3) spin rotational symmetry
with rotation angle $\theta\in[0,2\pi)$, and a similar form holds for $R$. Since $O_{\text{SO(3)}}^{\top}=O_{\text{SO(3)}}^{\dagger}$, the complex fermion operators transform as
@@ IV QMC simulations
In this section, we present details of our QMC simulations. We start by describing two different possible Trotter decompositions, demonstrate the absence of the sign problem, describe details of our algorithm, and eventually illustrate the convergence properties with respect to the limit of small Trotter time steps of the two different decompositions.
@@ IV.2 Absence of sign problem
IV.3 Algorithmic details
We use the the ALF implementation [39] of the auxiliary field QMC algorithm [40, 41, 42, 43]. This package makes use of the Gauss-Hermite quadrature to avoid sampling over continuous fields. Our finite inverse temperature $\beta$ calculations are carried out in the grand-canonical ensemble, choosing linear system sizes $L=6,9,12,15,18$ with periodic boundary conditions. For the zero-temperature results, we adopt a $\beta=L$ scaling, consistent with a dynamical critical exponent $z=1$, as naively expected from the linear fermion dispersion. If not stated otherwise, we use a Trotter time step of $\Delta\tau=0.2$. We choose units in which $t=1$ and $k_{\mathrm{B}}=1$, such that the temperature $T=1/\beta$ and the coupling constant $J\geq 0$ become dimensionless parameters, which we scan.
@@ V.1 GN-SO(3) transition
is the SO(3) spin structure factor at momentum $\boldsymbol{k}$ and imaginary time $\tau$, and $\boldsymbol{r}_{i}$ denotes the position vector of the lattice site $i$. The spin structure factor describes correlations between the spin density $\hat{\vec{s}}_{i,\lambda}(\tau)=\mathbf{\hat{c}}^{\dagger}_{i,\lambda}(\tau)\vec{K}\mathbf{\hat{c}}^{\phantom{\dagger}}_{i,\lambda}(\tau)$ (no summation over $\lambda$), where $\mathbf{\hat{c}}^{\dagger}_{i,\lambda}(\tau)$ is the time-evolved fermion operator in the Heisenberg picture. In order to locate the GN-SO(3) transition, we compute the renormalization group invariant SO(3) correlation ratio
@@
For increasing system sizes, the finite-size critical coupling approaches the thermodynamic critical point as $J_{\text{c1}}(L)=J_{\text{c1}}+aL^{-\omega-1/\nu}$, with nonuniversal coefficient $a$ [45]. Figure 5(a) shows the finite-size critical coupling as function of $1/L$ for two different size increments (green diamonds and purple squares) in comparison with the critical coupling obtained from the data-collapse analysis (orange dot). For sufficiently large lattice sizes, the estimates are consistent with each other within the numerical uncertainty.
@@
where $s(J,L)=\partial R_{\mathrm{c}}^{\mathrm{SO(3)}}(J,L)/\partial J$ corresponds to the slope of the correlation ratio as function of the coupling $J$. For increasing system sizes $L$, the deviation between the effective critical exponent and the corresponding value in the thermodynamic limit vanishes with $L^{-\omega}$. Our results for the effective finite-size critical exponents as function of $1/L$ for two different size increments (blue squares and orange circles) in comparison with the critical coupling obtained from the data-collapse analysis (green circles) in Figs. 5(b)–(d). For sufficiently large lattice sizes, the estimates from the different analyses are again consistent with each other within the numerical uncertainty.
Figure 5: (a) Finite-size critical coupling $J_{\text{c1}}(L)$ as function of $1/L$ from crossing-point analysis using size increments $c=3$ (purple squares) and $c=6$ (green diamonds), respectively, in comparison with the critical coupling $J_{\text{c1}}$ obtained from the data-collapse analysis (orange dot). For sufficiently large lattice sizes, the estimates are consistent with each other within the numerical uncertainty. (b) Same as (a), but for the effective correlation-length exponent $1/\nu(L)$. (c) Same as (a), but for the effective order-parameter anomalous dimension $\eta_{\phi}(L)$. (d) Same as (a), but for the effective fermion anomalous dimension $\eta_{\psi}(L)$. Blue dots in (b)–(d) show the field-theoretical estimates from Ref. [33] for comparison.
@@ V.2 SO(3)-U(1) transition
Figure 7: Spectral functions in (a) single-particle, (b) particle-hole, and (c) particle-particle channels at the SO(3)-U(1) transition point for $J\approx J_{\text{c2}}=1.057(10)$. Here, we show data for system size $L=12$ and inverse temperature $\beta=24$. Dash-dotted lines indicate the linear dispersions of the gapless excitations around the $\mathbf{K}$ (single-particle channel) and $\boldsymbol{\Gamma}$ (particle-hole and particle-particle channels) points in the Brillouin zone. The data support emergent Lorentz symmetry characterized by a single “velocity of light.”
@@
channels, where $|n\rangle$ and $E_{n}$, $n=0,1,2,\dots,$ correspond to the energy eigenstates and eigenvalues, respectively. Here, we consider $J\approx J_{\text{c2}}=1.057(10)$. As apparent from Fig. 7, all quantities show a linear dispersion relation at low energy. The single-particle spectral function is gapless at the $\mathbf{K}$ and $\mathbf{K}^{\prime}$ points in the Brillouin zone, while the particle-particle and particle-hole spectral functions are gapless at the $\boldsymbol{\Gamma}$ point. Importantly, the velocity in all three aforementioned channels are comparable, such that the data support emergent Lorentz symmetry characterized by a single “velocity of light.”
@@ VI Finite-temperature results
In the low-temperature phase for $T<T_{\mathrm{c}}^{\text{U(1)}}$, the system spontaneously selects an $s$-wave superconducting or interlayer-coherent insulating state as ground state. This implies that $\mathds{Z}_{2}$ PPH symmetry is broken in this phase as well. As a consequence, a lower bound for the critical temperature $T_{\mathrm{c}}^{\text{PPH}}$, marking the melting of PPH order, is given by the U(1) critical temperature $T_{\mathrm{c}}^{\text{U(1)}}$, i.e., $T_{\mathrm{c}}^{\text{PPH}}\geq T_{\mathrm{c}}^{\text{U(1)}}$. An interesting question is whether the two critical temperatures coincide or differ. The latter scenario would imply an intermediate vestigial phase for $T_{\mathrm{c}}^{\text{U(1)}}<T<T_{\mathrm{c}}^{\text{PPH}}$, in which PPH symmetry is spontaneously broken, but both U(1) global-charge and U(1) layer-charge symmetries remain intact. Such vestigial orders have previously been discussed in a variety of two-dimensional models [36, 37]. In order to characterize the melting of PPH order, we measure the PPH susceptibility
@@
where $\mathcal{S}_{\text{PPH}}$ is the PPH structure factor defined analogously to Eq. (44), using the parity operator $\hat{P}_{i}$ instead of the interlayer coherence operator $\hat{n}_{i}$. From the symmetry of the order parameter, we expect the PPH transition, if continuous, to fall into the 2D Ising universality class. In Fig. 9(c), we therefore show the rescaled PPH susceptibility $L^{\eta-2}\chi^{\mathrm{U(1)}}_{\boldsymbol{k}=\boldsymbol{\Gamma}}$ as function of $J$ for different system sizes $L$ and a representative fixed temperature [same temperature as those of the U(1) susceptibility shown in Fig. 9(a)], using the Ising exponent $\eta=0.25$. Similarly, we show in Fig. 9(d) the rescaled PPH susceptibility as function of $T$ for different system sizes $L$ and a representative fixed large value of $J$ [same coupling as those of the U(1) susceptibility shown in Fig. 9(b)] using $\eta=0.25$. While the data clearly indicate the PPH order at sufficiently large $J$ and low $T$, they are too noisy to clearly identify whether the critical temperature $T_{\mathrm{c}}^{\text{PPH}}$ lies above or right at $T_{\mathrm{c}}^{\text{U(1)}}$. This is also due to the fact that vestigial orders are typically realized in only small temperature windows above the corresponding primary orders [36, 37]. The presence of a tiny vestigial phase above $T_{\mathrm{c}}^{\text{U(1)}}$ can therefore at present not be excluded from our data. Exploring this possibility deserves further investigation.

### v3

# Critical properties of metallic and deconfined quantum phase transitions in Dirac systems

@@ I Introduction
In this work, we study a lattice model that features two metallic quantum phase transitions as function of coupling strength [32]. The first one is continuous and can be understood within the $(2+1)$-dimensional Gross-Neveu-SO(3) [GN-SO(3)] field theory [33]. In the SO(3)-broken phase, two out of three Dirac cones acquire a mass, and the wave function of the ungapped Dirac electron couples to the SO(3) order parameter. Similar mass terms have been put forward in spin-orbit-coupled fermions on a honeycomb lattice [34]. The second one is a transition between two long-range-ordered phases that break different symmetries: the aforementioned SO(3)-spin-symmetry-broken semimetal and a U(1)-charge-symmetry-broken fully gapped state  [^1] Note that the “fully gapped state” exhibits a finite gap in the single-particle spectrum while still retaining a gapless Goldstone excitation in the two-particle spectrum, associated with the spontaneously broken U(1) symmetry.. The latter can be understood as an insulator characterized by U(1) interlayer coherence [32]. We show, nevertheless, that the interlayer-coherent insulator is degenerate with an $s$-wave superconducting state as a consequence of a partial particle-hole (PPH) symmetry. At the level of mean-field theory, the transition between the SO(3)-spin- and U(1)-charge-symmetry-broken phases is strongly first order [32]. Quantum fluctuations beyond mean-field theory significantly weaken this order-to-order transition, rendering it a candidate for a deconfined quantum critical point in the presence of gapless fermionic excitations—a potential microscopic realization of metallic deconfined quantum criticality [36]. We use large-scale sign-problem free quantum Monte Carlo (QMC) simulations with a Hermitian Trotter decomposition, which significantly improves convergence properties with respect to the limit of small Trotter time steps, to compute the quantum critical and thermodynamic properties of both transitions. Furthermore, in contrast to our initial exploratory study of this model [32], we employ a microscopic implementation that preserves the model’s PPH symmetry explicitly, maintaining the degeneracy of the interlayer-coherent insulating and $s$-wave superconducting ground states in the fully gapped phase already on finite lattices. The interlayer-coherent insulating and $s$-wave superconducting states allow finite-temperature phases above the zero-temperature order. The latter breaks both U(1)-charge and PPH symmetries. Upon increasing temperature, it is therefore possible that the U(1) and PPH orders melt at different temperatures, with an intermediate vestigial phase at intermediate temperatures [37, 38, 39]. We compute the finite-temperature phase diagram of the model and show that the boundaries of the low-temperature orders vanish upon approaching the putative deconfined quantum critical point, in agreement with the expectation for a continuous or weak-first-order transition.
@@ II Model
Figure 1: (a) Sketch of bilayer honeycomb model defined by the Hamiltonian given in Eqs. (1)-(3). Each lattice site (purple dots) contain three SO(3) spin flavor degrees of freedom. The two layers interact via an SO(3)-symmetric spin-density interaction parametrized by the coupling $J$ (orange link). (c) Bottom: Schematic quantum phase diagram as function of $J$, showing the disordered Dirac semimetal (DSM) at small $J$, the SO(3)-spin-symmetry-broken Dirac semimetal (SO3 DSM), in which $2/3$ of the Dirac cones acquire a mass gap, while $1/3$ remains gapless, at intermediate $J$, as well as the fully gapped degenerate interlayer-coherent insulator (ILC) and $s$-wave superconductor (SSC) at strong $J$. The insets illustrate the fermion band structures in the three different phases. Top: Fermion single-particle spectral function from lattice mean-field theory along a high-symmetry path in the first Brillouin zone [depicted in (b)], illustrating the partial and full gap openings in the SO3 DSM and ILC/SSC phases, respectively.
@@ III.2 SO(3) spin rotational symmetry
with rotation angle $\theta\in[0,2\pi)$ and rotation axis $\vec{e}$, and a similar form holds for $R$. Since $O_{\text{SO(3)}}^{\top}=O_{\text{SO(3)}}^{\dagger}$, the complex fermion operators transform as
@@ IV QMC simulations
In this section, we present aspects of our QMC simulations. It is beyond the scope of this article to provide a detailed account of the algorithm and we will concentrate on model specific issues. We start by describing two different possible Trotter decompositions, demonstrate the absence of the sign problem, describe the specific implementation, and finally illustrate the convergence properties with respect to the limit of small Trotter time steps of the two different decompositions.
@@ IV.2 Absence of sign problem
IV.3 Specific implementation
We use the ALF implementation [41] of the auxiliary field QMC algorithm [42, 43, 44, 45]. This package utilizes Gauss-Hermite quadrature to replace sampling over continuous fields with a discrete field that takes the values ±$\pm 2$ and ±$\pm 1$. For further details, we refer to Refs. [41, 46]. While both approaches are formally equivalent, sampling over discrete fields generally reduces fluctuations. Our finite inverse temperature $\beta$ calculations are carried out in the grand-canonical ensemble, choosing linear system sizes $L=6,9,12,15,18$ with periodic boundary conditions. For the zero-temperature results, we adopt a $\beta=L$ scaling, consistent with a dynamical critical exponent $z=1$, as naively expected from the linear fermion dispersion. If not stated otherwise, we use a Trotter time step of $\Delta\tau=0.2$. We choose units in which $t=1$ and $k_{\mathrm{B}}=1$, such that the temperature $T=1/\beta$ and the coupling constant $J\geq 0$ become dimensionless parameters, which we scan.
@@ V.1 GN-SO(3) transition
is the SO(3) spin structure factor at momentum $\boldsymbol{k}$ and imaginary time $\tau$, and $\boldsymbol{r}_{i}$ denotes the position vector of the lattice site $i$. The spin structure factor describes correlations between the spin density $\hat{\vec{s}}_{i,\lambda}(\tau)=\mathbf{\hat{c}}^{\dagger}_{i,\lambda}(\tau)\vec{K}\mathbf{\hat{c}}^{\phantom{\dagger}}_{i,\lambda}(\tau)$ (no summation over $\lambda$), where $\mathbf{\hat{c}}^{\dagger}_{i,\lambda}(\tau)$ is the time-evolved fermion operator in the Heisenberg picture. In order to locate the GN-SO(3) transition, we compute the renormalization group invariant SO(3) correlation ratio [47]
@@
For increasing system sizes, the finite-size critical coupling approaches the thermodynamic critical point as $J_{\text{c1}}(L)=J_{\text{c1}}+aL^{-\omega-1/\nu}$, with nonuniversal coefficient $a$ [49]. Figure 5(a) shows the finite-size critical coupling as function of $1/L$ for two different size increments (orange diamonds and blue squares) in comparison with the critical coupling obtained from the data-collapse analysis (green circle). For sufficiently large lattice sizes, the estimates are consistent with each other within the numerical uncertainty.
@@
where $s(J,L)=\partial R_{\mathrm{c}}^{\mathrm{SO(3)}}(J,L)/\partial J$ corresponds to the slope of the correlation ratio as function of the coupling $J$. For increasing system sizes $L$, the deviation between the effective critical exponent and the corresponding value in the thermodynamic limit vanishes with $L^{-\omega}$. Our results for the effective finite-size critical exponents as function of $1/L$ for two different size increments (blue squares and orange diamonds) in comparison with the critical exponents obtained from the data-collapse analysis (green circles) are shown in Figs. 5(b)–(d). For sufficiently large lattice sizes, the estimates from the different analyses are again consistent with each other within the numerical uncertainty.
Figure 5: (a) Finite-size critical coupling $J_{\text{c1}}(L)$ as function of $1/L$ from crossing-point analysis using size increments $c=3$ (blue squares) and $c=6$ (orange diamonds), respectively, in comparison with the critical coupling $J_{\text{c1}}$ obtained from the data-collapse analysis (green dot). For sufficiently large lattice sizes, the estimates are consistent with each other within the numerical uncertainty. (b) Same as (a), but for the effective correlation-length exponent $1/\nu(L)$. (c) Same as (a), but for the effective order-parameter anomalous dimension $\eta_{\phi}(L)$. (d) Same as (a), but for the effective fermion anomalous dimension $\eta_{\psi}(L)$. Purple triangles in (b)–(d) show the field-theoretical estimates from Ref. [33] for comparison.
@@ V.2 SO(3)-U(1) transition
Figure 7: Spectral functions in (a) single-particle, (b) particle-hole, and (c) particle-particle channels at the SO(3)-U(1) transition point for $J=1.057\approx J_{\text{c2}}$. Here, we show data for system size $L=12$ and inverse temperature $\beta=24$. Dash-dotted lines indicate the linear dispersions of the gapless excitations around the $\mathbf{K}$ (single-particle channel) and $\boldsymbol{\Gamma}$ (particle-hole and particle-particle channels) points in the Brillouin zone. The data support emergent Lorentz symmetry characterized by a single “velocity of light.”
@@
channels, where $|n\rangle$ and $E_{n}$, $n=0,1,2,\dots,$ correspond to the energy eigenstates and eigenvalues, respectively. Here, we consider $J\approx J_{\text{c2}}=1.057(10)$. As apparent from Fig. 7, all quantities show a linear dispersion relation at low energy. The single-particle spectral function is gapless at the $\mathbf{K}$ and $\mathbf{K}^{\prime}$ points in the Brillouin zone, while the particle-particle and particle-hole spectral functions are gapless at the $\boldsymbol{\Gamma}$ point. The gapless excitations in the particle-particle and particle-hole channels can be understood as collective modes associated with the SO(3) and U(1) order parameters, which characterize the neighboring long-range-ordered phases. Importantly, the velocity in all three aforementioned channels are comparable, such that the data support emergent Lorentz symmetry characterized by a single “velocity of light.”
@@ VI Finite-temperature results
In the low-temperature phase for $T<T_{\mathrm{c}}^{\text{U(1)}}$, the system spontaneously selects an $s$-wave superconducting or interlayer-coherent insulating state as ground state. This implies that $\mathds{Z}_{2}$ PPH symmetry is broken in this phase as well. As a consequence, a lower bound for the critical temperature $T_{\mathrm{c}}^{\text{PPH}}$, marking the melting of PPH order, is given by the U(1) critical temperature $T_{\mathrm{c}}^{\text{U(1)}}$, i.e., $T_{\mathrm{c}}^{\text{PPH}}\geq T_{\mathrm{c}}^{\text{U(1)}}$. An interesting question is whether the two critical temperatures coincide or differ. The latter scenario would imply an intermediate vestigial phase for $T_{\mathrm{c}}^{\text{U(1)}}<T<T_{\mathrm{c}}^{\text{PPH}}$, in which PPH symmetry is spontaneously broken, but both U(1) global-charge and U(1) layer-charge symmetries remain intact. Such vestigial orders have previously been discussed in a variety of two-dimensional models [37, 38, 39]. In order to characterize the melting of PPH order, we measure the PPH susceptibility
@@
where $\mathcal{S}_{\text{PPH}}$ is the PPH structure factor defined analogously to Eq. (44), using the parity operator $\hat{P}_{i}$ instead of the interlayer coherence operator $\hat{n}_{i}$. From the symmetry of the order parameter, we expect the PPH transition, if continuous, to fall into the 2D Ising universality class. In Fig. 9(c), we therefore show the rescaled PPH susceptibility $L^{\eta-2}\chi^{\mathrm{U(1)}}_{\boldsymbol{k}=\boldsymbol{\Gamma}}$ as function of $J$ for different system sizes $L$ and a representative fixed temperature [same temperature as those of the U(1) susceptibility shown in Fig. 9(a)], using the Ising exponent $\eta=0.25$. Similarly, we show in Fig. 9(d) the rescaled PPH susceptibility as function of $T$ for different system sizes $L$ and a representative fixed large value of $J$ [same coupling as those of the U(1) susceptibility shown in Fig. 9(b)] using $\eta=0.25$. While the data clearly indicate the PPH order at sufficiently large $J$ and low $T$, they are too noisy to clearly identify whether the critical temperature $T_{\mathrm{c}}^{\text{PPH}}$ lies above or right at $T_{\mathrm{c}}^{\text{U(1)}}$. This is also due to the fact that vestigial orders are typically realized in only small temperature windows above the corresponding primary orders [37, 38, 39]. The presence of a tiny vestigial phase above $T_{\mathrm{c}}^{\text{U(1)}}$ can therefore at present not be excluded from our data. Exploring this possibility deserves further investigation.